OpenStudy (anonymous):

Prove 1/R = 1/R1 + 1/R2 + 1/R3 ?? (Resistant in parallel circuit)

OpenStudy (tcy):

\[I _{source}=I _{R1}+I_{R2}\] since Ohm's law: V=IR=> I=V/R \[\frac{ V _{source} }{ R _{T}}=\frac{ V _{R1} }{ R _{1}}+\frac{ V _{R2} }{ R _{2}}\] where VR1=VR2=Vsource for parallel

OpenStudy (anonymous):



OpenStudy (anonymous):

So in the dia, above,the current (I) is split into 3, siince I =V/R,the current is a third through each parallel lne.Hence resistance = 1/R for each leg